A railroad train travels forward along a straight Explanations for the answers please!
Answer #1

1.solution

You need to find out how long it takes to cover each 1000 m. The first segment takes (1000/80) seconds, the second takes (1000/50) seconds.

The average velocity is then the total distance 2000m divided by the sum of these two times.

(1,000/80) + (1,000/50) = 32.5 secs.
Distance = 2,000m.
(2,000/32.5) = average V of 61.53846m/sec

2.solution

2) By v = u - gt
=>0 = 19.6 - 9.8 x t
=>t = 2 sec
Thus total time of journey (T) = 2t = 4 sec
By v^2 = u^2 - 2gh
=>0 = (19.6)^2 - 2 x 9.8 x h
=>h = 19.6m
Thus total distance traveled (H) = 2h = 39.20m
Thus average speed = Total distance/Total time = 39.20/4 = 9.8 m/s